An angle measures 47 degrees . What is the measure of its complement?
Answer by fcabanski(1391)

The sum of complementary angles is 90 degrees.
47+x=90
x=90-47=43 degrees.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com